OpenStudy (anonymous):

write the slope-intercept form of the equation of each line 4x-3y=6

OpenStudy (hoblos):

slope-intercept form is of the form y=mx+b so 4x-3y=6 3y = 4x -6 y= (4/3)x -2

OpenStudy (anonymous):

thanks her is another one